# On the Completeness of Invariant Geometric Deep Learning Models

### v2

@@ F.1 Scaling to large graphs: MD22
AVG RANK 6.86 5.40 5.30 3.50 3.00 1.79 4.57 2.93
As presented in Table 3, we can see that GeoNGNN outperforms DisGNN across almost all targets. It also exhibits competitive results and achieves several new SOTA results on Stachyose and the two largest molecules, Buckyball catcher and Double-walled nanotube. On average, GeoNGNN performs better than all the equivariant models except for MACE, ranking the 2nd among all models.
We further evaluated the inference time and GPU memory consumption of GeoNGNN and several high-performance models. As shown in Table 4, GeoNGNN is more than 2x faster than MACE and more than 8x faster than Equiformer, which both utilize high-order equivariant representations and perform tensor products. Additionally, GeoNGNN demonstrates efficient memory usage, particularly when compared to 2-F-DisGNN and Equiformer. This could be attributed to the invariant representation and the simple architecture GeoNGNN adopts.
@@
Molecule # atoms MACE 2F-DisGNN Equiformer* GeoNGNN
Ac-Ala3-NHMe 42 0.166 0.130 0.305 0.070
13400 20542 38895 12997
Docosahexaenoic acid 56 0.209 0.215 0.397 0.088
17875 35844 51565 16980
Stachyose 87 0.388 OOM OOM 0.161
33146 OOM OOM 32430
AT-AT 60 0.195 0.244 0.361 0.083
17310 41125 50032 15864
AT-AT-CG-CG 118 0.437 OOM OOM 0.175
38090 OOM OOM 34951
Buckyball catcher 148 0.183 0.382 0.382 0.074
14466 62790 47702 13891
Double-walled nanotube 370 0.545 OOM OOM 0.180
44923 OOM OOM 36849
F.2 Generalization on out-of-domain data: 3BPA
GeoNGNN has demonstrated remarkable expressiveness. However, higher expressiveness only implies a better capability to fit data, and whether this leads to improved inductive bias and generalization power remains uncertain. Therefore, we conducted further evaluations of GeoNGNN on the 3BPA dataset [33], which includes lots of out-of-domain data in its test set, providing a good assessment of the generalization power of geometric models. The dataset comprises molecular dynamic simulations of the molecule 3-(benzyloxy)pyridin-2-amine, with the training set consisting of samples from the trajectory at 300K, and the test sets containing samples from trajectories at 300K, 600K, and 1200K.
@@ F.2 Generalization on out-of-domain data: 3BPA
Consistent with prior studies, we present the Root Mean Square Error for energy and force predictions, comparing GeoNGNN with MACE [3], NequIP [4], and Allegro [43]. The results are detailed in Table 5. GeoNGNN demonstrates competitiveness with Allegro, an equivariant model equipped with high-order tensors for learning local equivariant representations. However, when compared to the SOTA method MACE, GeoNGNN exhibits relatively worse performance, especially under higher temperatures, which signifies a greater distribution shift between the testing and training sets. This underscores that despite GeoNGNN’s strong theoretical expressiveness, its generalization power may lag behind that of those well-designed equivariant models. This discrepancy could potentially be attributed to the fact that GeoNGNN only adopts invariant representations, whereas equivariant representations could offer superior generalization abilities, particularly in sparse graph cases, as demonstrated in Du et al. [11]. Therefore, future research could focus on improving GeoNGNN’s generalization abilities by integrating additional equivariant representations, like the vectors used in [52, 54], or higher-order tensors applied in [3].
F.3 Effect of subgraph size
In this section, we delve into the impact of subgraph size, which serves as a form of receptive field for each node. Theoretically, a larger receptive field should always result in the same or better expressiveness. To explore how subgraph size affects experimental performance, we selected three molecules of different representative scales and observed the effects, as depicted in Figure 6. As anticipated, performance generally improves with increased subgraph size, roughly matching the improved theoretical expressiveness. However, the rule does not always hold: there are cases where increasing the subgraph size leads to degraded performance, which can be attributed to the potential noise and redundant information. It implies that there exists a conflict between theoretical expressiveness and experimental performance, and striking a balance between them is important.
@@ G.1 Dataset settings
We generate the synthetic dataset by implementing the counterexamples proposed by Li et al. [37].
The dataset includes 10 isolated cases and 7 combinatorial cases, where each case consists of a pair of symmetric point clouds that cannot be distinguished by DisGNN. Figure 7 shows two pairs of selected counterexamples.
Figure 7: Selected counterexamples in the synthetic dataset. Only colored nodes belong to the point clouds. The grey nodes and “edges" are for visualization purposes only. (a) One isolated case. (b) One combinatorial case.
@@
The 7 combinatorial cases are obtained through the augmentation of the 6 pairs of point clouds sampled from regular dodecahedrons and 1 pair from icosahedrons, using the method outlined in Theorem A.1 of Li et al. [37]. When augmenting the base point cloud, we consider the original, complementary, and all types of the base graph [37], respectively.
@@
The data split (training/validation/testing) in rMD17 is 950/50/the rest, following related works such as Batatia et al. [3], Gasteiger et al. [22]. For MD22, we adopt the data split specified in Chmiela et al. [9], which is also consistent with the other works. We train the models using a batch size of 4.
@@
$\mathcal{L}(\boldsymbol{X},\boldsymbol{z})=(1-\rho)|f_{\theta}(\boldsymbol{X},\boldsymbol{z})-\hat{t}(\boldsymbol{X},\boldsymbol{z})|+\frac{\rho}{N}\sum_{i=1}^{N}\sqrt{\sum_{\alpha=1}^{3}(-\frac{\partial f_{\theta}(\boldsymbol{X},\boldsymbol{z})}{\partial\boldsymbol{x}_{i\alpha}}-\hat{F}_{i\alpha}(\boldsymbol{X},\boldsymbol{z}))^{2}},$
where $X\in\mathbb{R}^{N\times 3}$ and $z\in\mathbb{R}^{N}$ represents the $n$ atoms’ coordinates and atomic number respectively, $\hat{t}$ and $\hat{F}$ represents the molecule’s energy and forces acting on atoms respectively. The force ratio $\rho$ is chosen as 0.99 or 0.999.
In rMD17, the experimental results of NequIP [4], GemNet [22], are sourced from Musaelian et al. [43]. The result of PaiNN [52] is sourced from Batatia et al. [3]. The results of 2-F-DisGNN [37] and MACE [3] are reported by their original papers. We trained and evaluated DimeNet [20] using its implementation in Pytorch Geometric [16] since the original paper did not report results on this dataset, and we use the default parameters (We find that the default parameters produce the best performance).
In MD22, the results of sGDML and MACE [34] are sourced from Kovacs et al. [34]. The results of PaiNN [52], TorchMD-Net [54], Allegro [43], and Equiformer [38] are obtained from the work presented by Li et al. [36].
@@
Following Batatia et al. [3],we split the training set and validation set at a ratio of 450/50, utilizing pre-split test sets for evaluation. We use a batch size of 4, an identical objective function to that in rMD17 and MD22, and a force ratio of 0.999 were employed. The results of all other models are sourced from Batatia et al. [3].
G.2 Model settings.
@@
Following previous work, we use radial basis functions (RBF) $f^{\rm rbf}_{\rm e}:\mathbb{R}\to\mathbb{R}^{H_{\rm rbf}}$ to expand the euclidean distance between two nodes into a vector, which is shown to be beneficial for inductive bias Gasteiger et al. [20], Li et al. [37]. We use the same expnorm RBF function as Li et al. [37], defined as
@@
We set the number of VD-Conv layers to 7 for rMD17 and 6 for MD22. The hidden dimension is set to 512 and the dimension of radial basis functions (RBF) is set to 16. The message passing cutoff is set to $13$Å for rMD17 and $7$Å for MD22. We employ the polynomial envelope with $p=6$, as proposed in Gasteiger et al. [20], along with the corresponding cutoff.
@@
The high-level architecture of GeoNGNN is described in Definition 4.1. The inner DisGNN, which runs on the ego subgraph of each node, is similar to the DisGNN block shown in Figure 8. However, the inner DisGNN additionally performs node marking and distance encoding for each node $j$ in node $i$’s subgraph before passing their representations to VD-Conv layers:
@@
The hidden dimension and message passing cutoff are set to the same values as those in DisGNN. The subgraph size is set to 13 Å for rMD17, 5 Å for MD22 and 6 Å for 3BPA, and the maximal subgraph size is set to 25. We set $(N_{\text{in}},N_{\text{out}})$ as $(5,2)$ for rMD17, $(3,3)$ for MD22 and $(6,2)$ for 3BPA. It is worth mentioning that the total number of convolutional layers is the same as that of DisGNN. This ensures that the receptive field of the two models is roughly the same, thereby guaranteeing a relatively fair comparison.
We optimize all models using the Adam optimizer [32], incorporating exponential decay and plateau decay learning rate schedulers, as well as a linear learning rate warm-up. To mitigate the risk of overfitting, we employ early stopping based on validation loss and apply exponential moving average (EMA) with a decay rate of 0.99 to the model parameters during the validation and testing phases. The models are trained on Nvidia RTX 4090 and Nvidia A100 (80GB) GPUs. For rMD17, training hours for each molecule range from 20 to 70. For MD22, the training hours ranged from 35 to 150 on Nvidia 4090 for all molecules except Stachyose (due to longer convergence time) and Double-walled nanotube (the largest molecule necessitating an A100), which required about 210 and 70 GPU hours on the Nvidia 4090 and 80GB A100, respectively. For 3BPA, the training hours is around 10 GPU hours on Nvidia 4090.
@@ H.1 Definition and theoretical analysis of GeoNGNN-C
After three rounds of the Chiral DisGNN (i.e., the DisGNN that incorporates the directed distance as a geometric representation as described in Definition H.1), node representations $h_{j}^{(3)}$ for any node $j$ can encode the directed distances $d_{ji}$, $d_{jc}$, and $d_{ic}$. This enables the completion of the triangular distance encoding necessary for reconstruction in Lemma E.6.
Next, we employ a modified version of the proof from Lemma E.6 to demonstrate that with an additional two rounds of the Chiral DisGNN, we can derive the SE(3)- and permutation-invariant identifier of the point cloud $P$.
In the proof of Lemma E.6, we first search across all the nodes to find two nodes $a$ and $b$ that form the minimal dihedral angles $\theta(aic,bic)$, and then identifies the coordinates of $a,b,i,c$ up to Euclidean isometry from the all-pair distances $D=\{(m,n,d_{mn})\mid m,n\in\{a,b,i,c\}\}$. However, since Chiral DisGNN embeds the geometry using directed distance, we now have $D_{i}=\{(m,n,d_{i,mn})\mid m,n\in\{a,b,i,c\}\}$ instead. We are now able to identify the coordinates of $a,b,i,c$ up to SE(3)-transformation (i.e., Euclidean isometry without reflection) from $D_{i}$ in the following way:
@@
Note that the orientation relevant to SE(3) transformation is already determined at this stage. We proceed with the remaining steps of the proof outlined in Lemma E.6 and reconstruct the whole geometry in a deterministic way based on the known 4-tuple $abic$, and therefore reconstruct the whole geometry up to permutation and SE(3)-transformation.

### v3

@@ F.2 Scaling to Large Graphs: MD22
[…]
In this table, we further include a recent SOTA model VisNet-LSRM (Li et al., 2023). As shown in Table 3, GeoNGNN consistently outperforms DisGNN across all targets. It also delivers competitive results compared to other SOTA models, including MACE and VisNet-LSRM, ranking 2nd on average. These results demonstrate the effectiveness of subgraph enhancement in practical GeoNGNN with finite configurations.
To explore how subgraph size affects experimental performance, we selected three molecules of different representative scales and observed the effects, as depicted in Figure 6. As anticipated, performance generally improves with increased subgraph size, roughly matching the improved theoretical expressiveness. However, the rule does not always hold: there are cases where increasing the subgraph size leads to degraded performance, which can be attributed to the potential noise and redundant information. It implies that practically, there could exist a conflict between theoretical expressiveness and experimental performance, and striking a balance between them is important.
We further evaluated the inference time and GPU memory consumption of GeoNGNN and several high-performance models. As shown in Table 4, GeoNGNN is more than 2x faster than MACE and more than 8x faster than Equiformer, which both utilize high-order equivariant representations and perform tensor products. Additionally, GeoNGNN demonstrates efficient memory usage, particularly when compared to Equiformer. This could be attributed to the invariant representation and the simple architecture GeoNGNN adopts.
@@
Molecule # atoms MACE Equiformer* GeoNGNN
Ac-Ala3-NHMe 42 0.166 0.305 0.070
13400 38895 12997
Docosahexaenoic acid 56 0.209 0.397 0.088
17875 51565 16980
Stachyose 87 0.388 OOM 0.161
33146 OOM 32430
AT-AT 60 0.195 0.361 0.083
17310 50032 15864
AT-AT-CG-CG 118 0.437 OOM 0.175
38090 OOM 34951
Buckyball catcher 148 0.183 0.382 0.074
14466 47702 13891
Double-walled nanotube 370 0.545 OOM 0.180
44923 OOM 36849
F.3 Generalization on Out-of-Domain Data: 3BPA
GeoNGNN has demonstrated remarkable expressiveness. However, higher expressiveness only implies a better capability to fit data, and whether this leads to improved inductive bias and generalization power remains uncertain. Therefore, we conducted further evaluations of GeoNGNN on the 3BPA dataset (Kovács et al., 2021), which includes lots of out-of-domain data in its test set, providing a good assessment of the generalization power of geometric models. The dataset comprises molecular dynamic simulations of the molecule 3-(benzyloxy)pyridin-2-amine, with the training set consisting of samples from the trajectory at 300K, and the test sets containing samples from trajectories at 300K, 600K, and 1200K.
@@ F.3 Generalization on Out-of-Domain Data: 3BPA
Consistent with prior studies, we present the Root Mean Square Error for energy and force predictions, comparing GeoNGNN with MACE (Batatia et al., 2022), NequIP (Batzner et al., 2022), and Allegro (Musaelian et al., 2023). The results are detailed in Table 5. GeoNGNN demonstrates competitiveness with Allegro, an equivariant model equipped with high-order tensors for learning local equivariant representations. However, when compared to the SOTA method MACE, GeoNGNN exhibits relatively worse performance, especially under higher temperatures, which signifies a greater distribution shift between the testing and training sets. This underscores that despite GeoNGNN’s strong theoretical expressiveness, its generalization power may lag behind that of those well-designed equivariant models. As stated in the main text Section 6, this discrepancy could potentially be attributed to the fact that GeoNGNN only adopts invariant representations, whereas equivariant representations could offer superior generalization abilities, particularly in sparse graph cases, as demonstrated in Du et al. (2024).
F.4 Handling Various Property Predictions: QM9
Finally, we evaluate GeoNGNN on QM9 (Ramakrishnan et al., 2014; Wu et al., 2018), which contains approximately 13k well-annotated small molecules, each paired with 12 properties to predict. The results, shown in Table 6, indicate that GeoNGNN performs particularly well in energy-related predictions, such as $U_{0}$ and $H$.
[…]
Target Unit SchNet PhysNet ComENet SphereNet 2F-Dis. PaiNN Torchmd DisGNN GeoNGNN
[…]
AVG RANK 8.42 7.25 4.58 3.42 3.67 3.92 3.17 7.42 2.92
@@ G.1 Dataset Settings
We generate the synthetic dataset by implementing the counterexamples proposed by Li et al. (2024).
The dataset includes 10 isolated cases and 7 combinatorial cases, where each case consists of a pair of symmetric point clouds that cannot be distinguished by DisGNN. Figure 7 shows two pairs of selected counterexamples, and more counterexamples are provided in Figure 8.
Figure 7: Selected counterexamples in the synthetic dataset. Only colored nodes belong to the point clouds. The grey nodes and “edges” are for visualization purposes only. (a) One isolated case. (b) One combinatorial case.
Figure 8: Pairs of point clouds that cannot be distinguished by DisGNN, taken from Li et al. (2024). The nodes are arranged on the surfaces of regular polyhedra, only red or blue nodes are part of point clouds; the grey nodes and the “edges” are included solely for vi- sualization purposes. For those subfigures with two node colors, nodes of the same color in the left and right point clouds represent one specific pair of indistinguishable point clouds by GeoNGNN. The label "Scale cube" signifies that the relative sizes of the two regular polyhedra can vary arbitrarily. Each subfigure, except the bottom-rightmost one, represents an isolated counterexample. The bottom-rightmost case illustrates a combinatorial counterexample, constructed by combining multiple instances of the top-left pair in an "origin-all-complementary" pattern. Comprehensive explanations of these counterexamples are provided in Li et al. (2024).
@@
The 7 combinatorial cases are obtained through the augmentation of the 6 pairs of point clouds sampled from regular dodecahedrons and 1 pair from icosahedrons, using the method outlined in Theorem A.1 of Li et al. (2024). When augmenting the base point cloud, we consider the original, complementary, and all types of the base graph (Li et al., 2024), respectively.
@@
The data split (training/validation/testing) in rMD17 is 950/50/the rest, following related works such as Batatia et al. (2022); Gasteiger et al. (2021). For MD22, we adopt the data split specified in Chmiela et al. (2023), which is also consistent with the other works. We train the models using a batch size of 4.
@@
$\mathcal{L}(\bm{X},\bm{z})=(1-\rho)|f_{\theta}(\bm{X},\bm{z})-\hat{t}(\bm{X},\bm{z})|+\frac{\rho}{m}\sum_{i=1}^{n}\sqrt{\sum_{\alpha=1}^{3}(-\frac{\partial f_{\theta}(\bm{X},\bm{z})}{\partial\bm{x}_{i\alpha}}-\hat{F}_{i\alpha}(\bm{X},\bm{z}))^{2}},$
where $X\in\mathbb{R}^{n\times 3}$ and $z\in\mathbb{R}$ represents the $n$ atoms’ coordinates and atomic number respectively, $\hat{t}$ and $\hat{F}$ represents the molecule’s energy and forces acting on atoms respectively. The force ratio $\rho$ is chosen as 0.99 or 0.999.
In rMD17, the experimental results of NequIP (Batzner et al., 2022), GemNet (Gasteiger et al., 2021), are sourced from Musaelian et al. (2023). The result of PaiNN (Schütt et al., 2021) is sourced from Batatia et al. (2022). The results of 2-F-DisGNN (Li et al., 2024) and MACE (Batatia et al., 2022) are reported by their original papers. We trained and evaluated DimeNet (Gasteiger et al., 2019) using its implementation in Pytorch Geometric (Fey & Lenssen, 2019) since the original paper did not report results on this dataset, and we use the default parameters (We find that the default parameters produce the best performance).
In MD22, the results of sGDML and MACE (Kovacs et al., 2023) are sourced from Kovacs et al. (2023). The results of PaiNN (Schütt et al., 2021), TorchMD-Net (Thölke & De Fabritiis, 2021), Allegro (Musaelian et al., 2023), and Equiformer (Liao & Smidt, 2022) are obtained from the work presented by Li et al. (2023).
@@
Following Batatia et al. (2022), we split the training set and validation set at a ratio of 450/50, utilizing pre-split test sets for evaluation. We use a batch size of 4, an identical objective function to that in rMD17 and MD22, and a force ratio of 0.999 were employed. The results of all other models are sourced from Batatia et al. (2022).
QM9
Following Li et al. (2024), we split the training set and validation set at a ratio of 110K/10K, utilizing pre-split test sets for evaluation. We use a batch size of 32. See the following for details of model hyper-parameters.
G.2 Model Settings
@@
Following previous work, we use radial basis functions (RBF) $f^{\rm rbf}_{\rm e}:\mathbb{R}\to\mathbb{R}^{H_{\rm rbf}}$ to expand the euclidean distance between two nodes into a vector, which is shown to be beneficial for inductive bias Gasteiger et al. (2019); Li et al. (2024). We use the same expnorm RBF function as Li et al. (2024), defined as
@@
We set the number of VD-Conv layers to 7 for rMD17 and 6 for MD22. The hidden dimension is set to 512 and the dimension of radial basis functions (RBF) is set to 16. The message passing cutoff is set to $13$Å for rMD17 and $7$Å for MD22. We employ the polynomial envelope with $p=6$, as proposed in Gasteiger et al. (2019), along with the corresponding cutoff.
NGNN
GeoNGNN is based on the high-level framework of NGNN proposed in (Zhang & Li, 2021). NGNN learns over topological graphs by nesting a base GNN such as GCN (Kipf & Welling, 2016), GIN (Xu et al., 2018b), or GraphSAGE (Hamilton et al., 2017), as illustrated in Figure 10. For completeness, we include formal descriptions provided in (Zhang & Li, 2021).
Formally, given a topological graph $G=(V,E)$, where $V=\{1,2,\dots,n\}$ represents the set of nodes and $E\subseteq V\times V$ represents the set of edges, NGNN first defines $k$-hop ego subgraphs $G_{k}^{v}$ for each node $v\in V$. Each $k$-hop ego subgraph $G_{k}^{v}$ is induced by the nodes within $k$-hops from $v$ (including $v$ itself) and the edges connecting these nodes.
For each rooted subgraph $G_{k}^{w}$, NGNN applies a base GNN to perform $T^{\text{in}}$-rounds of message passing. Let $v$ be any node appearing in $G_{k}^{w}$. Denote the hidden state of $v$ at time $t$ in subgraph $G_{k}^{w}$ as $h_{v,G_{k}^{w}}^{t}$. The initial hidden state $h_{v,G_{k}^{w}}^{0}$ is typically set to the node’s raw features, such as embeddings of atomic numbers. When explicit node marking is adopted, which is the default in GeoNGNN, $h_{v,G_{k}^{w}}^{0}$ is further fused with a unique mark embedding to distinguish the root node within its subgraph.
After $T^{\text{in}}$-rounds of message passing, a subgraph pooling operation aggregates the node embeddings $\{\!\!\{h_{v,G_{k}^{w}}^{T^{\text{in}}}\mid v\in G_{k}^{w}\}\!\!\}$ into a single subgraph representation $h_{G_{k}^{w}}$:
$h_{G_{k}^{w}}=R_{\text{subgraph}}\left(\{\!\!\{h_{v,G_{k}^{w}}^{T^{\text{in}}}\mid v\in G_{k}^{w}\}\!\!\}\right),$
where $R_{\text{subgraph}}$ is the subgraph pooling function that summarizes all node-level information within the subgraph. This subgraph representation $h_{G_{k}^{w}}$ serves as the final representation of the root node $w$ in the original graph.
Subsequently, an outer GNN performs $T^{\text{out}}$-rounds of message passing, obtaining the hidden representation $h_{v}^{t}$ of node $v$ at step $t$, where the initial representation is $h_{v}^{0}=h_{G_{k}^{v}}$. Finally, the graph-level representation $h_{G}$ is produced:
$h_{G}=R_{\text{graph}}\left(\{\!\!\{h_{v}^{T^{\text{out}}}\mid v\in V\}\!\!\}\right),$
where $R_{\text{graph}}$ is a global pooling function that aggregates node-level embeddings into a comprehensive graph-level representation. This process enables NGNN to capture hierarchical representations of the graph, leveraging both local (subgraph-level) and global (graph-level) features.
Figure 10: Illustration of NGNN (Zhang & Li, 2021), taken from  (Zhang & Li, 2021). NGNN operates by first extracting rooted subgraphs around each node in the original graph. A base GNN with a subgraph pooling layer is then applied independently to each rooted subgraph to compute its representation. The resulting subgraph representation serves as the original feature for the root node in the original graph. Subsequently, an outer GNN is applied to the original graph, leveraging these node features and original graph structures to extract final graph-level representations.
@@
GeoNGNN serves as the geometric counterpart to NGNN, operating on point clouds rather than topological graphs in Figure 10. The high-level architecture of GeoNGNN is described in the main body. The inner DisGNN, which runs on the ego subgraph of each node, is similar to the DisGNN block shown in Figure 9. However, the inner DisGNN additionally performs node marking and distance encoding for each node $j$ in node $i$’s subgraph before passing their representations to VD-Conv layers:
@@
The hidden dimension and message passing cutoff are set to the same values as those in DisGNN. The subgraph size is set to 13 Å for rMD17, 5 Å for MD22, 6 Å for 3BPA, 13 Å for QM9 and the maximal subgraph size is set to 25. We set $(N_{\text{in}},N_{\text{out}})$ as $(5,2)$ for rMD17, $(3,3)$ for MD22, $(6,2)$ for 3BPA and $(4,1)$ for QM9. It is worth mentioning that the total number of convolutional layers is the same as that of DisGNN. This ensures that the receptive field of the two models is roughly the same, thereby guaranteeing a relatively fair comparison.
We optimize all models using the Adam optimizer (Kingma & Ba, 2014), incorporating exponential decay and plateau decay learning rate schedulers, as well as a linear learning rate warm-up. To mitigate the risk of overfitting, we employ early stopping based on validation loss and apply exponential moving average (EMA) with a decay rate of 0.99 to the model parameters during the validation and testing phases. The models are trained on Nvidia RTX 4090 and Nvidia A100 (80GB) GPUs. For rMD17, training hours for each molecule range from 20 to 70. For MD22, the training hours ranged from 35 to 150 on Nvidia 4090 for all molecules except Stachyose (due to longer convergence time) and Double-walled nanotube (the largest molecule necessitating an A100), which required about 210 and 70 GPU hours on the Nvidia 4090 and 80GB A100, respectively. For 3BPA, the training hours is around 10 GPU hours on Nvidia 4090. For QM9, the training hour is around 36 to 100 on the Nvidia 4090.
@@ H.1 Definition and Theoretical Analysis of GeoNGNN-C
After three rounds of the Chiral DisGNN (i.e., the DisGNN that incorporates the directed distance as a geometric representation as described in Definition H.1), node representations $h_{j}^{(3)}$ for any node $j$ can encode the directed distances $d_{ji}$, $d_{jc}$, and $d_{ic}$. This enables the completion of the triangular distance encoding necessary for reconstruction in Lemma B.6.
Next, we employ a modified version of the proof from Lemma B.6 to demonstrate that with an additional two rounds of the Chiral DisGNN, we can derive the SE(3)- and permutation-invariant identifier of the point cloud $P$.
In the proof of Lemma B.6, we first search across all the nodes to find two nodes $a$ and $b$ that form the minimal dihedral angles $\theta(aic,bic)$, and then identifies the coordinates of $a,b,i,c$ up to Euclidean isometry from the all-pair distances $D=\{(m,n,d_{mn})\mid m,n\in\{a,b,i,c\}\}$. However, since Chiral DisGNN embeds the geometry using directed distance, we now have $D_{i}=\{(m,n,d_{i,mn})\mid m,n\in\{a,b,i,c\}\}$ instead. We are now able to identify the coordinates of $a,b,i,c$ up to SE(3)-transformation (i.e., Euclidean isometry without reflection) from $D_{i}$ in the following way:
@@
Note that the orientation relevant to SE(3) transformation is already determined at this stage. We proceed with the remaining steps of the proof outlined in Lemma B.6 and reconstruct the whole geometry in a deterministic way based on the known 4-tuple $abic$, and therefore reconstruct the whole geometry up to permutation and SE(3)-transformation.